40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1867     Entry time: Thu Aug 17 19:15:00 2017
Author: Craig, awade 
Type: DailyProgress 
Category: BEAT 
Subject: Strong precavity beatnote found near DC, both cavities locked 

We locked both cavities with a -4.67 dBm beatnote at 140 MHz.

This elog is so we remember the settings of everything.  Notably, the South Slow Control Voltage is -6.7333 V and the North Slow Control Voltage is 2.6925 V.

The South path slow PID loop seems to be out of tune, and is causing the PZT to rail and eventually we lose the lock.  Perhaps negative voltages were not anticipated by the creator of the slow control PID locker, or something out of the range of -2 to 7 volts, which were the orignal rails until awade changed them last week.  We will have to look at this.  For now we shut down the South PID and are manually tuning the temp so the SCAV_FAST_MON monitoring the PZT voltage is not too great.

For now, we are hooking up the transmission transimpedence PD in hopes of getting a transmission beatnote.  We are adjusting the North cavity temp to bring the 140 MHz beatnote pictured down closer to DC.  The beatnote is already at 110 MHz as I write this.  The North cavity temperature control power supply reads 10.5 V, whereas before it was 12 V.

 

Attachment 1: medmFSS.jpg  3.650 MB  | Hide | Hide all
medmFSS.jpg
Attachment 2: lockedCavities.jpg  2.078 MB  | Hide | Hide all
lockedCavities.jpg
Attachment 3: beatnote145MHz.jpg  1.655 MB  | Hide | Hide all
beatnote145MHz.jpg
ELOG V3.1.3-